OpenStudy (anonymous):

Bob is driving through town. Two traffic radar detectors are spaced 0.75 miles apart. He passes the first and is clocked at 30 mph. One minute later, he passes the second and is clocked at 34mph. Show mathematically that somewhere between the two that he exceeded the 35mph speed limit. Fully justify your answer, citing precise facts.

OpenStudy (anonymous):

This is can be proved using the mean value theorem for integral. There exist a time t between (0,1) so such that f(t) (1) = 0.75 f(t) = 0.75 mi/min converting to mile/hr and you get 45 mi/hr. This mean at some time between 0 and 1min, his speed equals to the average speed.

OpenStudy (anonymous):

so the fact that his speed was 30 mi/hr at t = 0 and 40 mi/hr at t = 1 is irrelevant.

OpenStudy (anonymous):

but wasnt he at 34mi/hr at t = 1?

OpenStudy (anonymous):

that doesn't mean he didn't speed up sometime between t = 0min and t = 1min, which is garanteed by the mean value theorem for integral

OpenStudy (anonymous):

okay so then f'(c) = f(1)-f(0)/1 = 45mph?

OpenStudy (anonymous):

what do you mean by f'(c) ?

OpenStudy (anonymous):

I am just trying to relate what you said to what I have in my notes. I have: Suppost f(x) is continuous on [a, b], and f(x) if differentiable on (a, b). Then there is at least one number c in (a, b) where the average slope = the exact slope formula: f'(c)= f(b)-f(a)/b-a

OpenStudy (anonymous):

ok, that works. The important thing is you need understand what f represents. f(1) - (0) is the change in distance which is 0.75 miles. And a = 0 , b = 1, 0.75 / (1- 0) = 0.75 mi/min (which is 45 mi/hr) So the the Mean Value theorem, there is a c in (0,1) such that f'(c) = 0.75 mi/hr

OpenStudy (anonymous):

Now do you see why his initial speed and final speed recorded by the speed camera are irrelevant? The average speed is 45 mi/hr. We just need to show at some point his speed equals the average speed 45, which is greater than speed limit

OpenStudy (anonymous):

Yes thanks you but just clarifying just in case for f'(c) you meant .75 mi/min in the very last line of your response.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I don't get what you just said

OpenStudy (anonymous):

in your response relating your answer to my notes at the end you said f'(c) = 0.75 mi/hr . I am just wondering if that was a typo and it should be f'(c) = 0.75 mi/min? or is it suppose to be as you put it?

OpenStudy (anonymous):

ah yes. That was my typo. Should be 0.75 miles/min

OpenStudy (anonymous):

Alright thank you very much!

OpenStudy (anonymous):

no prob :)